OpenStudy (anonymous):

81^(2x)=9 , how can I find x?

OpenStudy (anonymous):

4x=1

OpenStudy (anonymous):

x=1/4 lols

OpenStudy (anonymous):

Write \(81\) as \(9^2\). So, \(9^{4x}=9\). Since they have the same base, they must have the same exponent. So \(4x=1\), which gives \(x= \frac{1}{4}\).

OpenStudy (anonymous):

81 = 9^2 , so 81^(2x) = (9^2)^(2x) = 9^(4x) = 9 , 4x=1 etc

OpenStudy (anonymous):

oh, gotcha. Than kyou :)